# Accelerated Rotation-Invariant Convolution for UAV Image Segmentation

### v1

@@ Abstract
Rotation invariance is essential for precise, object-level segmentation in UAV aerial imagery, where targets can have arbitrary orientations and exhibit fine-scale details. Conventional segmentation architectures like U-Net rely on convolution operators that are not rotation-invariant, leading to degraded segmentation accuracy across varying viewpoints. Rotation invariance can be achieved by expanding the filter bank across multiple orientations, however, this will significantly increase computational cost and memory traffic. In this paper, we introduce a GPU-optimized rotation-invariant convolution framework that eliminates the traditional data-lowering (im2col) step required for matrix-multiplication-based convolution. By exploiting structured data sharing among symmetrically rotated filters, our method achieves multi-orientation convolution with greatly reduced memory traffic and computational redundancy. We further generalize the approach to accelerate convolution with arbitrary (non-symmetric) rotation angles. Across extensive benchmarks, the proposed convolution achieves 20–55% faster training and 15–45% lower energy consumption than cuDNN, while maintaining accuracy comparable to state-of-the-art rotation invariant methods.. In the eight-orientation setting, our approach achieves up to 45% speedup and 41% energy savings on 256×256 inputs, and 32% speedup and 23% lower energy usage on 1024×1024 inputs. Integrated into a U-Net segmentation model, the framework yields up to 6% improvement in accuracy over the non-rotation-aware baseline. These results demonstrate that the proposed method provides an effective and highly efficient alternative to existing rotation invariant CNN frameworks.
@@ I Introduction
Robust semantic segmentation in high‑resolution aerial images requires rotation invariance, because objects of interest—whether vegetation, buildings, vehicles, or other land‑surface features—can occur at any orientation and often contain complex fine‑scale detail. Yet the convolution operator used in mainstream deep learning segmentation architectures (e.g., U‑Net) is translation equivariant but not rotation invariant, leaving performance sensitive to object orientation. Convolution is fundamental in deep learning due to its ability to efficiently extract spatial and hierarchical features from input data [cnn-survey]. Convolution layers routinely account for the bulk of total FLOPs, memory traffic, and energy use in convolution‑based networks. They therefore dominate training time and bound inference throughput, making their optimization a central focus of prior works [FFT1], [winograd], [10286398]. These costs are further amplified when pursuing rotation‑invariant convolution, which typically requires evaluating kernels across multiple orientation samples. Despite the additional computational burden, achieving rotation invariance is essential for many aerial remote sensing applications—including satellite imagery [satellite2, satellite3, satellite4], UAV imaging [UAV1], and geospatial image localization [geo1]—because feature extraction must remain consistent under arbitrary rotations of the target object in the image; this need is further underscored by recent UAV studies in geology (dual-spectrum hot-spring fluid segmentation) [YI2025661], urban scene segmentation (UAVformer) [YI2023109019], and UAV visual perception (CCTseg) [YI2023112612] under varying altitudes and viewpoints.
@@
The extracted feature maps are rotation equivariant, i.e., feature maps whose outputs are rotated in a predictable way, as we rotate the input. A special case of equivariance is invariance [harmonic-cnn], where the output values of the feature map are not affected by rotations of the input. In essence, the rotation invariance can be achieved by directly encoding rotation equivariance into the convolution layer followed by a global pooling operation. Multiple rotated versions of the canonical filter are applied at the same input to generate rotation equivariant feature maps. For instance, Cohen & Welling [G] propose a group convolution, code-name G-convolution, in which symmetric rotations (rotation angle is multiple of 90 degrees) are applied to the same filter, generating a rotation equivariant feature map. However, such rotation equivariant convolution introduces significant overhead on the computation. For example, with G-convolution, the computational overhead increases by 4 times compared to conventional convolution in 2D space, since four symmetric rotations are applied at each filter. A single rotation equivariant convolution is equivalent to four conventional convolutions. The computational overhead possibly increases by 24 times in 3D space because there are 24 symmetric transformations available in 3D space [CubeNet1]. The number of symmetric rotations and therefore the computational overhead increases along with increasing dimension of the input data. In this paper, we show how to use the proposed convolution operation to accelerate this kind of rotation equivariant convolution which is based on the filter weight sharing scheme.
@@
As described in our previous work [manduhu2025airbornesensedetectdrones], conventional (gather‑style) convolution accumulates contributions at each output location: every neighboring input pixel is multiplied by its corresponding kernel weight, and the results are summed into the central output position. In contrast, our scatter convolution reverses this data flow: each input pixel is multiplied by a bank of filter weights and the resulting products are scattered directly to their destination output locations. We originally introduced this formulation to support sparse convolutions on irregularly sampled LiDAR data. In this paper, we generalize the scatter approach to (i) dense, standard grid convolution and (ii) rotation‑invariant convolution. The scatter mapping removes the im2col‑style data‑lowering step typically required by matrix multiplication based implementations, reducing memory traffic and simplifying the kernel. Moreover, because the scatter mapping exposes weight sharing across symmetric rotations, intermediate results can be reused—collapsing the number of distinct multiplications to one per symmetry group and substantially lowering computational cost.
@@
  • Comprehensive evaluation: Experimental results demonstrate substantial throughput gains and energy savings across all evaluated workloads, with the proposed method achieving approximately 20–55% faster training and 15–45% lower energy consumption compared with cuDNN. The approach maintains accuracy competitive with state-of-the-art rotation-equivariant methods, and when integrated into a baseline network without rotation handling, it further improves segmentation accuracy by up to 6%.
@@ II-A Segmentation of remote sensing imagery
Transformer-based architectures have also been used for aerial image segmentation. SegHSI [10751785] introduces a head-free backbone with cluster attention and a spatial-aware feed-forward network to capture spectral–spatial correlations in HSI data. The work in [10422983] employs a transformer based encoder–decoder for UAV forestry segmentation, enabling accurate delineation of diseased pine canopies in high resolution imagery.
@@ III Methodology
III-A Acceleration of Single-Orientation Convolution
@@
III-B Acceleration of symmetric rotation invariant convolution
@@
Here $R_{-r}$ denotes a rotation of the gradient by $r\times 90^{\circ}$, mapping each rotated kernel’s gradient back to the base orientation.
@@ III-C Parallelization of the scatter operation
III-D Acceleration of arbitrary rotation convolution
@@ IV-A GPU implementation of single-orientation convolution with scatter operation
Figure 8: Output image tiling scheme for scattering operation with on-chip memory of GPU, where the size of overlap between tiles is $\lfloor k/2\rfloor$.
@@ V-A Model for Evaluation
To evaluate the performance of the scatter based rotation invariant convolution in a real segmentation application, we replaced the first convolution of base convolution block in UNet [ronneberger2015u] with the rotation invariant convolution described above (see Fig. 9 for illustration). To balance performance and the memory-expansion problem, we apply max pooling across every four symmetric rotations; for example, with eight rotations, this produces two orientation features, each corresponding to a group of four symmetric rotations. The second convolution within the base block uses a standard convolution, we found through experiments that this design improves performance.
@@ V-B2 Semantic Segmentation Drone dataset
The Semantic Drone Segmentation dataset contains 400 high-resolution nadir-view images of urban residential scenes, captured at altitudes between 5 m and 30 m above ground and annotated with detailed pixel-level semantic labels covering 24 object categories. In our experiments, we use 328 images for training, 32 for validation, and 40 for testing.
@@ V-C Evaluation with Different Convolution Approaches
V-C1 Single orientation convolution
@@
We compare our approach with the state-of-the-art ORN [orn] and E2CNN [e2cnn] frameworks, both of which provide APIs for achieving rotation invariant. Tables I and II present the performance of cuDNN, ORN, E2CNN, and our scatter-based convolution under the single-orientation setting. Because single-orientation rotation convolution degenerates to standard convolution, all methods achieve nearly identical accuracy. Nevertheless, meaningful differences appear in computational cost. On the plant dataset with 256×256 inputs, scatter-based convolution trains approximately 9.2% faster than cuDNN and reduces energy consumption by about 6.9%. At 1024×1024 resolution, scatter achieves roughly a 7.0% reduction in training time and a 7.6% reduction in energy usage. Similar trends are seen on the drone dataset, where scatter-based convolution is about 9.8% faster at 256 resolution and 7.1% faster at 1024 resolution, while saving approximately 5.9% and 7.9% energy, respectively. These results show that even in the absence of multi-orientation modeling, the scatter based convolution offers consistent efficiency gains without sacrificing accuracy. Fig. 10 and Fig. 11 show representative segmentation predictions for plant dataset and drone view datasets.
@@
V-C2 Symetric four orientation convolution
Tables III and IV summarize the results for the symmetric four-orientation configuration. Relative to single-orientation convolution, accuracy improves slightly, particularly for E2CNN, reflecting the benefit of limited rotation modeling. The computational advantages of scatter-based convolution become more pronounced under this setting. On the plant dataset, scatter reduces training time by approximately 45.8% at 256 resolution and 28.8% at 1024 resolution, while lowering energy consumption by about 38.4% and 18.5%, respectively. On the drone dataset, scatter provides comparable benefits, achieving about 46.7% faster training at 256 resolution and 32.9% at 1024 resolution, with corresponding energy reductions of roughly 39.1% and 22.8%. These results indicate that symmetric four-orientation convolutions provide a strong balance between improved segmentation accuracy and computational efficiency, with scatter based convolution offering the best performance–efficiency trade-off.
@@ V-C2 Symetric four orientation convolution
V-C3 Symetric eight orientation convolution
Tables V and VI report the eight-orientation results, which represent the highest degree of rotation invariance in our experiments. Increasing the number of modeled orientations leads to further accuracy improvements, especially for the plant dataset. The scatter-based convolution continues to provide substantial efficiency benefits. On the plant dataset, scatter is approximately 45.4% faster than cuDNN at 256 resolution and about 20.8% faster at 1024 resolution, while reducing energy usage by approximately 41.2% and 12.9%, respectively. On the drone dataset, scatter achieves around 56.9% faster training at 256 resolution and 32.9% at 1024 resolution, with energy savings of about 44.3% and 22.8%. These comparisons show that as rotation resolution increases, rotation-invariant convolutions become more accurate, and the scatter-based implementation remains significantly more efficient than ORN and E2CNN, consistently delivering between 20% and 55% faster training and 15% to 45% lower energy consumption.
@@ V-C3 Symetric eight orientation convolution
1024 cudnn based conv 81361.0 5.553 75.87
ORN conv 79179.0 5.412 73.64
E2CNN conv 73003.0 5.024 75.64
scatter based conv 64434.0 4.835 75.85
@@
1024 cudnn based conv 30936.2 2.103 92.44
ORN conv 29956.3 2.056 91.20
E2CNN conv 27548.6 1.894 91.99
scatter based conv 20761.8 1.623 92.40

### v2

@@ Abstract
Rotation invariance is essential for precise object level segmentation in UAV aerial imagery, where targets can have arbitrary orientations and exhibit fine scale details. Conventional segmentation architectures like UNet rely on convolution operators that are not rotation invariant, leading to degraded segmentation accuracy across varying viewpoints. Rotation invariance can be achieved by expanding the filter bank across multiple orientations; however, this significantly increases computational cost and memory requirement. In this paper, we introduce a GPU optimized rotation invariant convolution framework that eliminates the traditional data lowering (im2col) step required for matrix multiplication based convolution. By exploiting structured data sharing among symmetrically rotated filters, our method achieves multi orientation convolution with greatly reduced memory requirements and computational redundancy. We further generalize the approach to accelerate convolution with arbitrary (non symmetric) rotation angles. Integrated into a UNet segmentation model, the framework yields up to a 5.7% improvement in accuracy over the non rotation aware baseline. Across extensive benchmarks, the proposed convolution achieves 20–57% faster training and 15–45% lower energy consumption than cuDNN, while maintaining accuracy comparable to state-of-the-art rotation invariant methods. Because the scatter based operator greatly reduces intermediate feature dimensionality, the efficiency of our design also enables practical sixteen orientation convolution and pooling, yielding further accuracy gains that are infeasible for conventional rotation invariant implementations. These results demonstrate that the proposed method provides an effective and highly efficient alternative to existing rotation invariant convolution frameworks.
@@ I Introduction
Robust semantic segmentation in high‑resolution aerial images requires rotation invariance, because objects of interest—whether vegetation, buildings, vehicles, or other land‑surface features—can occur at any orientation and often contain complex fine‑scale detail. Yet the convolution operator used in mainstream deep learning segmentation architectures (e.g., U‑Net) is translation equivariant but not rotation invariant, leaving performance sensitive to object orientation. Convolution is fundamental in deep learning due to its ability to efficiently extract spatial and hierarchical features from input data [cnn-survey]. Convolution layers routinely account for the bulk of total FLOPs, memory traffic, and energy use in convolution‑based networks. They therefore dominate training time and bound inference throughput, making their optimization a central focus of prior works [FFT1], [winograd], [10286398].
These costs are further amplified when pursuing rotation‑invariant convolution, which typically requires evaluating kernels across multiple orientation samples. Despite the additional computational burden, achieving rotation invariance is essential for many aerial remote sensing applications—including satellite imagery [satellite2, satellite3, satellite4], UAV imaging [UAV1], and geospatial image localization [geo1]—because feature extraction must remain consistent under arbitrary transformations of the target object in the image, including rotations; this need is further underscored by recent UAV studies in geology (dual-spectrum hot-spring fluid segmentation) [YI2025661], urban scene segmentation (UAVformer) [YI2023109019], and UAV visual perception (CCTseg) [YI2023112612] under varying altitudes and viewpoints. The method in [mitton2021rotation] embeds rotation equivariance into U-Net for UAV image segmentation, achieving a 7% improvement in deforestation segmentation accuracy.
@@
The extracted feature maps are rotation equivariant, i.e., feature maps whose outputs are rotated in a predictable way, as we rotate the input. A special case of equivariance is invariance [harmonic-cnn], where the output values of the feature map are not affected by rotations of the input. In essence, the rotation invariance can be achieved by directly encoding rotation equivariance into the convolution layer followed by a global pooling operation. Multiple rotated versions of the canonical filter are applied at the same input to generate rotation equivariant feature maps. For instance, Cohen & Welling [G] propose a group convolution, code-name G-convolution, in which symmetric rotations (rotation angle is multiple of 90 degrees) are applied to the same filter, generating a rotation equivariant feature map. However, such rotation equivariant convolution introduces significant overhead on the computation. For example, with G-convolution, the computational overhead increases by 4 times compared to conventional convolution in 2D space, since four symmetric rotations are applied at each filter. The computational overhead possibly increases by 24 times in 3D space because there are 24 symmetric transformations available in 3D space [CubeNet1]. The number of symmetric rotations and therefore the computational overhead increases along with increasing dimension of the input data. As reported in prior works [11125636, yang2025group], the increased number of symmetric rotations leads to substantial additional feature computations, making the training of rotation-invariant convolutional networks significantly slower than that of conventional convolutional models. In this paper, we show how the proposed convolution operation can be used to accelerate rotation invariant convolution by eliminating the data lowering step and exploiting filter weight sharing.
@@
As described in the previous work [manduhu2025airbornesensedetectdrones], conventional (gather‑style) convolution accumulates contributions at each output location: every neighboring input pixel is multiplied by its corresponding kernel weight, and the results are summed into the central output position. In contrast, our scatter convolution reverses this data flow: each input pixel is multiplied by a bank of filter weights and the resulting products are scattered directly to their destination output locations. We originally introduced this formulation to support sparse convolutions on irregularly sampled LiDAR data. In this paper, we generalize the scatter approach to (i) dense, standard grid convolution and (ii) rotation‑invariant convolution. The scatter mapping removes the im2col‑style data‑lowering step typically required by matrix multiplication based implementations, reducing memory traffic and simplifying the kernel. Moreover, because the scatter mapping exposes weight sharing across symmetric rotations, intermediate results can be reused—collapsing the number of distinct multiplications to one per symmetry group and substantially lowering computational cost.
While computational efficiency is an important consideration, achieving rotation invariance in UAV image segmentation inherently involves a trade-off between segmentation accuracy and computational efficiency. Increasing the number of modeled orientations generally improves accuracy by enhancing rotational robustness and feature expressiveness; however, this improvement comes at the cost of increased computation, memory traffic, and intermediate feature-map growth. As demonstrated in our experiments, moderate orientation resolutions (e.g., eight orientations) provide a favorable balance by significantly improving segmentation accuracy while remaining computationally feasible. In contrast, further increasing the orientation resolution leads to prohibitively large intermediate feature maps for conventional rotation-invariant implementations, making them infeasible to execute within GPU memory constraints. This work therefore focuses not only on improving computational efficiency, but also on enabling practical high-resolution rotation handling by reducing redundant computation and memory overhead, allowing finer orientation sampling under realistic GPU constraints.
@@
  • Comprehensive evaluation: Integrated into a UNet model, the proposed convolution yields up to a 5.7% improvement over the non rotation aware baseline, while achieving 20–57% faster training and 15–45% lower energy consumption than cuDNN across all evaluated workloads. Its efficient scatter based design also enables practical sixteen orientation convolution, providing additional gains unattainable with conventional rotation invariant methods.
@@ II-A Segmentation of remote sensing imagery
Attention-based architectures have also been used for aerial image segmentation. SegHSI [10751785] adopts a CNN-based backbone augmented with cluster attention to model spectral–spatial correlations in hyperspectral imagery. The work in [10422983] employs a Transformer–CNN hybrid encoder–decoder for UAV forestry segmentation, enabling accurate delineation of diseased pine canopies in high resolution imagery.
@@ III Methodology
UAV aerial image segmentation presents unique challenges compared with ground-level imagery, as target objects often appear at arbitrary orientations due to the aerial imaging perspective. Conventional convolution operators are sensitive to such orientation variations, which can lead to degraded segmentation accuracy. The methodology presented in this section introduces a series of convolution formulations that are specifically designed as drop-in replacements for standard convolution layers in semantic segmentation networks.These operators form the computational backbone of the proposed UAV segmentation framework, enabling rotation-aware feature extraction while maintaining efficiency suitable for large-scale aerial imagery. Their practical effectiveness is demonstrated through integration into a U-Net architecture and validated in the experimental section.
III-A Single orientation convolution formulation
@@
III-B Symmetric rotation-invariant convolution formulation
@@
Here $R_{-r}$ denotes a rotation of the gradient by $r\times 90^{\circ}$, mapping each rotated kernel’s gradient back to the base orientation. The proposed scatter approach can also accelerate backpropagation, since the gradient computations themselves are convolution operations, as shown in these formulations.
@@ III-C Parallelization of the scatter operation
III-D Arbitrary rotation convolution formulation
@@ IV-A GPU implementation of single-orientation convolution with scatter operation
Figure 8: Image tiling with on-chip memory of GPU, where the size of overlap between tiles is $\lfloor k/2\rfloor$.
@@ V-A Model for Evaluation
To evaluate the performance of the scatter based rotation invariant convolution in a real segmentation application, we replaced the first convolution of base convolution block in UNet [ronneberger2015u] with the rotation invariant convolution described above (see Fig. 9 for illustration). To balance performance and dimension explosion, we apply max pooling across every four symmetric rotations; for example, with eight rotations, this produces two orientation features, each corresponding to a group of four symmetric rotations. The second convolution within the base block uses a standard convolution, we found through experiments that this design improves performance.
@@ V-B2 Semantic Segmentation Drone dataset
The Semantic Drone Segmentation dataset contains 400 high-resolution nadir-view images of urban residential scenes, captured at altitudes between 5 m and 30 m above ground and annotated with detailed pixel-level semantic labels covering 24 object categories. In our experiments, we use 340 images for training, 32 for validation, and 28 for testing.
@@ V-C Evaluation with Different Convolution Approaches
V-C1 Single orientation rotation convolution
@@
We compare our approach with cuDNN as well as the state-of-the-art ORN [orn] and E2CNN [e2cnn] frameworks, both of which provide APIs for achieving rotation invariance. Tables I and II report the performance of cuDNN, ORN, E2CNN, and the proposed scatter-based convolution under the single-orientation setting. The observed accuracy variation is within $\pm 0.1\%$. Since single-orientation rotation convolution degenerates to standard convolution, all methods achieve nearly identical segmentation accuracy. Nevertheless, notable differences arise in computational efficiency. On the plant dataset with 256×256 inputs, E2CNN achieves the lowest training time and energy consumption, while scatter-based convolution also demonstrates competitive efficiency, training approximately 9.2% faster than cuDNN and reducing energy consumption by about 6.9%. At 1024×1024 resolution, scatter-based convolution achieves the best efficiency, reducing training time by roughly 7.0% and energy usage by 7.6% compared with cuDNN. Similar trends are observed on the drone dataset: at 256 resolution, E2CNN attains the highest accuracy and lowest computational cost, whereas scatter-based convolution remains consistently more efficient than cuDNN. At 1024 resolution, scatter-based convolution outperforms all baselines in both training time (7.1% reduction) and energy consumption (7.9% reduction). These results indicate that while E2CNN is particularly efficient for smaller input resolutions, the proposed scatter-based convolution provides consistent and scalable efficiency gains—especially at higher resolutions—without sacrificing segmentation accuracy, as it avoids explicit data lowering and reduces memory traffic through direct scatter accumulation. Representative segmentation results are visulized in Fig. 10 and Fig. 11.
@@
V-C2 Symetric four orientation rotation convolution
Tables III and IV summarize the results for the symmetric four-orientation configuration. Relative to single-orientation convolution, accuracy increases by about 2%, with E2CNN achieving the highest accuracy, reflecting the benefit of four rotation modeling. The computational advantages of scatter-based convolution become more pronounced under this setting. On the plant dataset, scatter reduces training time by approximately 45.9% at 256 resolution and 28.8% at 1024 resolution, while lowering energy consumption by about 38.4% and 18.6%, respectively. On the drone dataset, scatter provides comparable benefits, achieving about 46.7% faster training at 256 resolution and 32.9% at 1024 resolution, with corresponding energy reductions of roughly 39.1% and 22.8%. These results indicate that symmetric four-orientation convolutions achieve strong segmentation accuracy with manageable computational cost. The scatter based convolution provides the most favorable accuracy–efficiency profile, benefiting from reuse of intermediate computations across symmetric rotations. This enables substantial reductions in training time and energy consumption while maintaining segmentation accuracy.
@@ V-C2 Symetric four orientation rotation convolution
V-C3 Symetric eight orientation rotation convolution
Tables V and VI report the eight-orientation results. Increasing the number of modeled orientations leads to further accuracy improvements, especially for the plant dataset. Because convolutional pooling over eight orientations produces twice the feature-map dimensionality of the four-orientation configuration. However, it imposes substantially greater memory and computational demands on conventional implementations. The scatter based convolution continues to provide substantial efficiency benefits. On the plant dataset, scatter is approximately 45.4% faster than cuDNN at 256 resolution and about 34.5% faster at 1024 resolution, while reducing energy usage by approximately 41.2% and 21.7%, respectively. On the drone dataset, scatter achieves around 56.9% faster training at 256 resolution and 32.7% faster at 1024 resolution, with energy savings of about 44.3% and 22.7%, respectively.
The proposed scatter-based approach differs from E2CNN in how rotation is modeled. While E2CNN learns orientation-dependent expansion coefficients, the scatter-based method constructs steerable filters from a set of learnable base filters, as described in Section III-D, where orientation dependence is introduced analytically. By enforcing orthogonality and equal-magnitude constraints on these base filters, the formulation approximates rotation invariance, leading to accuracy trends distinct from E2CNN.
The experimental results demonstrate that as rotation resolution increases, rotation invariant convolutions become more accurate, and the scatter based implementation remains more efficient than ORN and E2CNN. ORN shows smaller accuracy improvements in the eight-orientation setting because it does not provide an API for orientation-subgroup pooling; in our implementation, its group pooling operation aggregates all eight orientations into a single pooled feature. As accuracy continues to improve, conventional rotation-invariant convolutions incur rapidly increasing memory and computation overhead due to expanded feature-map dimensionality. This effect is particularly evident for E2CNN, where group representations lead to larger intermediate tensors. By contrast, the proposed scatter-based implementation mitigates this issue by consolidating rotation handling within a single convolution pass, thereby maintaining scalability as orientation resolution increases.
@@ V-C3 Symetric eight orientation rotation convolution
1024 cudnn based conv 100269.79 6.8162 75.86
ORN conv 95240.30 6.5967 74.19
E2CNN conv 84227.06 5.8908 75.57
scatter based conv 65703.51 5.3363 75.82
@@
1024 cudnn based conv 41836.42 2.8440 92.44
ORN conv 41382.05 2.8402 91.20
E2CNN conv 36342.18 2.4986 91.99
scatter based conv 28137.53 2.1996 92.40
V-C4 Symetric sixteen orientation rotation convolution
In preliminary experiments, we found that modeling sixteen orientations in UNet leads to very large intermediate feature tensors and impractical memory and compute overhead. Compared to implementations that process each rotation separately, our scatter-based approach performs symmetric four-rotation convolution and pooling in a single pass within GPU on-chip memory, reducing the number of intermediate feature maps by a factor of four. As a result, our scatter-based implementation is the only approach capable of supporting a full sixteen-orientation convolution and pooling within practical memory and runtime constraint. As shown in Tables VII and VIII, the scatter-based approach exhibits improved accuracy when employing sixteen orientations relative to the eight-orientation configuration. Because convolutional pooling over sixteen orientations generates twice the feature-map dimensionality of the eight-orientation configuration.
TABLE VII: Plant Data Segmentation Using Sixtee-Orientation Rotation-Invariant Convolution
256 scatter based conv 31208.3 1.555 79.27
1024 scatter based conv 103612.2 6.685 76.33
TABLE VIII: Semantic drone segmentation Using Sixteen-Orientation Rotation-Invariant Convolution
256 scatter based conv 9982.3 0.513 94.39
1024 scatter based conv 41905.6 3.101 92.87